OpenStudy (anonymous):

HOW DOES AN INCREASE OF 1 IN THE INDEPENDENT VARIABLE X AFFECT f(x)? f(x)=x^3 A. f(x) will have a change of 3x^2+3x+1 B. f(x) will have a change of x^3-3x+1 C. f(x) will have a change of x^3+3x+1 D. f(x) will have a change of 1-3x^2

OpenStudy (anonymous):

I am thinking it is C? Is that right?

OpenStudy (loser66):

nope

OpenStudy (anonymous):

Then it's B

OpenStudy (anonymous):

I dont think it is A or D

OpenStudy (loser66):

hahaha... you GGGGGGGGGGGGGUess

OpenStudy (loser66):

It's A

OpenStudy (anonymous):

WHAT!!?!?! ITS NOT B??

OpenStudy (anonymous):

HOW?!?!?

OpenStudy (loser66):

\(x \rightarrow x+1\) \(x^3 \rightarrow (x+1)^3 = x^3 +3x^2+3x+1\) hence the changing is \(x^3 +3x^2+3x+1\color {red}{-} x^3 = A\)

OpenStudy (loser66):

ok??

OpenStudy (anonymous):

Omg yes! hahaha how silly of me!

OpenStudy (anonymous):

Thank you!

OpenStudy (loser66):

@gabymarie313 I can't find your post, tag me there